(5,1) is a solution t the equation 2x-3y=7.what does this mean?

Answer by Maths68(1474)

2x-3y=7
If we have above equation we don't know what is value of x and y.
But a solution has given which (5,1). It means that (x,y)=(5,1)
x=5
y=7
When we put these values in given equation we have
2x-3y=7
2(5)-3(1)=7
10-3=7
7=7
It means that (5,1) is a solution of above equation